If a 10 gram sample of a radioactive substance has a half-life of 6000 years, how much would be present after 8000 years?
Answer #1

Ans :

Half life formula is given as :

Nt = No (1/2)t / h

Nt is amount of sample left

No is initial amount of sample = 10 g

t = time lapsed = 8000 y

h = half life = 6000 y

Nt = 10 g (1/2)8000 y / 6000 y

= 3.97 g

The amount of sample left will be 3.97 g.
Answer #2

Step 1: Understand the Half-Life Formula

The amount of substance remaining after a given time can be calculated using the formula:

N=N0×(12)tt1/2

where:

  • N = remaining quantity after time t,

  • N0 = initial quantity,

  • t = elapsed time,

  • t1/2 = half-life of the substance.

Step 2: Plug in the Given Values

  • Initial quantity, N0=10 grams,

  • Half-life, t1/2=6000 years,

  • Elapsed time, t=8000 years.

N=10×(12)80006000

Step 3: Simplify the Exponent

80006000=431.3333

Step 4: Calculate the Remaining Amount

N=10×(12)1.3333N10×0.3969N3.969 grams

Answer is :-

After 8000 years, approximately 3.97 grams of the radioactive substance would remain


answered by: Harshwardhan kunal